A sequence {mn} is recursively defined as
m1=2 mn+1= 3/5-mn
Write the first four terms of this sequence

Answer by Theo(13342)

it looks like they're using M for the term, rather than A which is normally used.

i read this as:

m(1) = 2
m(n+1) = 3/5 - m(n)

n represents the term number which starts at 1.
m represents the value of each term.
m(1) is the value of the first term.
m(2) is the value of the second term.
m(n) is the value of the nth term.
m(n+1) is the value of the term right after the nth term.

you start with m(1).
m(n+1) is then equal to m(2).
m(n+1) is then equal to m(3).
m(n+1) is then equal to m(4).
etc.....

your recursive formula is m(n+1) = 3/5 - m(n).

so your sequence starts as follows:

m(1) = 2
m(2) = 3/5 - 2 = 3/5 - 10/5 = -7/5.
m(3) = 3/5 - (-7/5) = 3/5 + 7/5 = 10/5
m(4) = 3/5 - 10/5 = -7.5
m(5) = 3/5 - (-7.5) = 3/5 + 7/5 = 10/5
etc.

it looks like this recursive formula is going to switch back and forth from 2 to -7/5 to 2 to -7/5 forever.

since 2 is equal to 10/5, this is the same as saying that this recursive formula is going to switch back and forth from 10/5 to -7/5 to 10/5 to -7/5 forever.

the first 4 terms are:

m(1) = 2
m(2) = -7/5
m(3) = 2
m(4) = -7/5